OpenStudy (anonymous):

Assuming x ≠ 0 and y ≠ 0, what is the quotient of 8x^9y^3 +2x^6 y^2+8x^3y/2x^3y

OpenStudy (anonymous):

Long division!

OpenStudy (anonymous):

divide term by term

OpenStudy (anonymous):

wat did u do @jiteshmeghwal9

OpenStudy (anonymous):

\[\frac{8x^9y^3}{2x^3y}=4x^{9-3}y^{3-1}=4x^6y^2\] for exampole then move on to the next one and the next one

OpenStudy (jiteshmeghwal9):

sorry ! i think that x=0 & y=0

OpenStudy (anonymous):

4x^3y^3 + x^2y^2 + 4xy??

OpenStudy (anonymous):

second term should have \(x^3\)

OpenStudy (anonymous):

don't you divide the exponents? thats where i get lost.

OpenStudy (anonymous):

oh i see you subtract the exponents not divide

OpenStudy (anonymous):

so is 4x^6y^2 + x^3y + 4 the answer @satellite73

OpenStudy (anonymous):

it is really cancelling for example \[\frac{x^6}{x^3}=\frac{x\times x\times x\times x\times x\times x}{x\times x\times x}\] \[=\frac{\cancel{x}\cancel{x}\cancel{x}x\times x\times x}{\cancel{x}\cancel{x}\cancel{x}}=x^3\]

OpenStudy (anonymous):

yes you got it

OpenStudy (anonymous):

thanks